### v1

# Recoloring via modular decomposition

Abstract
The reconfiguration graph of the $k$-colorings of a graph $G$, denoted $R_{k}(G)$, is the graph whose vertices are the $k$-colorings of $G$ and two colorings are adjacent in $R_{k}(G)$ if they differ in color on exactly one vertex. A graph $G$ is said to be recolorable if $R_{\ell}(G)$ is connected for all $\ell\geq\chi(G)$+1. We use the modular decomposition of several graph classes to prove that the graphs in the class are recolorable. In particular, we prove that every ($P_{5}$, diamond)-free graph, every ($P_{5}$, house, bull)-free graph, and every ($P_{5}$, $C_{5}$, co-fork)-free graph is recolorable.
1 Introduction
Let $G$ be a finite simple graph with vertex-set $V(G)$ and edge-set $E(G)$. We assume that $G$ is connected unless stated otherwise. We use $n$ to represent the number of vertices in a graph. Two vertices $u$ and $v$ are adjacent in $G$ if $uv\in E(G)$. For a positive integer $k$, a $k$-coloring of $G$ is a mapping from $V(G)$ to a set of colors $\{1,2,\dots,k\}$ such that no two adjacent vertices receive the same color. We say that $G$ is $k$-colorable if it admits a $k$-coloring, and the chromatic number of $G$, denoted $\chi(G)$, is the minimum number of colors required to color $G$. We say $\chi$-coloring of $G$ instead of $\chi(G)$-coloring of $G$, when appropriate. The reconfiguration graph of the $k$-colorings, denoted $R_{k}(G)$, is the graph whose vertices are the $k$-colorings of $G$ and two colorings are adjacent in $R_{k}(G)$ if they differ in color on exactly one vertex. A graph $G$ is said to be $k$-mixing if $R_{k}(G)$ is connected. Given a reconfiguration graph, we can ask the following questions: Is the reconfiguration graph connected? If so, what is the diameter of the reconfiguration graph? In this paper we focus on the connectivity of the reconfiguration graph of the $k$-colorings. We say a graph $G$ is recolorable if $R_{\ell}(G)$ is connected for all $\ell\geq\chi(G)$+1.
Deciding whether there exists a path between two colorings in $R_{k}(G)$ is PSPACE-complete for $k>3$ [6], and can be decided in polynomial time for $k\leq 3$ [8]. In 2018, Wrochna proved that the problem remains PSPACE-complete for graph classes with bounded bandwidth, and hence for graph classes with bounded treewidth [18]. This led researchers to study the problem for various restricted graph classes. In [5], a graph class known as $k$-color-dense graphs was introduced and it was proved that every $k$-color-dense graph is recolorable. The class of $k$-color-dense graphs includes the class of $k$-colorable chordal graphs. For a graph $H$, a graph $G$ is $H$-$free$ if no induced subgraph of $G$ is isomorphic to $H$. In [2], it was proved that every $H$-free graph $G$ is recolorable if and only if $H$ is an induced subgraph of $P_{4}$ or $P_{3}$+$P_{1}$. For a collection of graphs $\mathcal{H}$, $G$ is $\mathcal{H}$-free if $G$ is $H$-free for every $H\in\mathcal{H}$.
Bonamy and Bousquet [4] used the tree decomposition of graphs to show that every graph $G$ is ($tw(G)$+2)-mixing, where $tw(G)$ denotes the treewidth of the graph $G$. In this paper we use the modular decomposition of a graph $G$ to prove that $G$ is recolorable.
A set $S\subseteq V(G)$ is a module if $1\leq|S|\leq|V(G)|$ and every vertex in $V(G)\setminus S$ is either adjacent to every vertex of $S$ or no vertex of $S$. A module $S$ is said to be non-trivial if $S$ contains at least two vertices. Note that for each $v\in V(G)$, $\{v\}$ is a trivial module in $G$. A graph $G$ is prime if it does not contain any non-trivial module.
By substituting a graph $H$ for $S\subseteq V(G)$ we mean taking the graph $G$-$S$ and adding an edge between every vertex of $H$ and every vertex of $G$-$S$ that is adjacent to some vertex of $S$ in $G$. Note that this is slightly different from the well-known definition of substituting a graph for a vertex. However, substituting a graph $H$ for $S\subseteq V(G)$ is equivalent to contracting the set $S$ to a vertex $v$ (and removing any loops and multiple edges created), and then applying the usual definition of substituting the graph $H$ for the vertex $v$. A blowup of a graph $G$ is a graph obtained by substituting non-empty cliques for some vertices of $G$.
(a) $diamond$
(b) $house$
(c) $bull$
(d) $co$-$fork$
Figure 1:
A class of graphs $\mathcal{G}$ is said to be hereditary if for every $G$ in $\mathcal{G}$, every induced subgraph of $G$ is also in $\mathcal{G}$. Note that for a collection of graphs $\mathcal{H}$, the class of $\mathcal{H}$-free graphs is hereditary. In Section 3, we prove the following.
Theorem 1.
Let $\mathcal{G}$ be a hereditary class of graphs and let $\mathcal{H}$ be the set of prime graphs in $\mathcal{G}$. If for all $H\in\mathcal{H}$, every blowup of $H$ is recolorable, then for all $G\in\mathcal{G}$, $G$ is recolorable.
We use the above result in Section 4 along with the modular decomposition of several subclasses of $P_{5}$-free graphs to prove the following.
Theorem 2.
Every ($P_{5}$, diamond)-free graph is recolorable.
Theorem 3.
Every ($P_{5}$, house, bull)-free graph is recolorable.
Note that there exist $P_{5}$-free graphs that are not recolorable [11]. A graph $G$ is said to be $P_{4}$-sparse if every set of five vertices in $G$ induces at most one $P_{4}$. The class of $P_{4}$-sparse graphs generalizes the class of $P_{4}$-free graphs. Jamison and Olariu [16] gave a forbidden induced subgraph characterization for $P_{4}$-sparse graphs: A graph $G$ is $P_{4}$-sparse if and only if $G$ is ($P_{5}$, $C_{5}$, house, fork, co-fork, banner, co-banner)-free. In 1997, Fouquet and Giakoumakis [13] introduced the class of semi-$P_{4}$-sparse graphs which is a generalization of the class of $P_{4}$-sparse graphs. A graph $G$ is semi-$P_{4}$-sparse if $G$ is ($P_{5}$, $C_{5}$, co-fork)-free. We prove the following.
Theorem 4.
Every semi-$P_{4}$-sparse graph is recolorable.
Feghali and Fiala [10], proved that every 3-colorable ($P_{5}$, house, $C_{5}$)-free graph is recolorable and asked if every ($P_{5}$, house, $C_{5}$)-free graph is recolorable. We prove the following.
Theorem 5.
Every ($P_{5}$, house)-free graph is recolorable if and only if every ($P_{5}$, house, $C_{5}$)-free graph is recolorable.
Section 5 contains some open questions.
2 Preliminaries
A path in a graph is a sequence of distinct vertices $v_{i}$ and edges of the form $v_{0}$, $v_{0}v_{1}$, $v_{1},\dots,v_{p-1}$, $v_{p-1}v_{p}$, $v_{p}$. The length of a path is equal to the number of edges in the path. A graph is said to be connected if there exists a path between every pair of distinct vertices of the graph. A vertex $v\in V(G)$ is said to be universal if $v$ is adjacent to every vertex in $V(G)\setminus\{v\}$. We say a coloring $\alpha$ of $G$ is reachable from a coloring $\beta$ of $G$ in $R_{\ell}(G)$ if there is a path between $\alpha$ and $\beta$ in $R_{\ell}(G)$. Where $A\subseteq V(G)$ and $B\subseteq V(G)$, we say $A$ is complete to $B$ if every vertex in $A$ is adjacent to every vertex in $B$. A set of mutually adjacent vertices in a graph is called a clique. A set of mutually non-adjacent vertices in a graph is called an independent set. Where $\alpha$ is a coloring of $G$ and $S\subseteq V(G)$, we use $\alpha(S)$ to denote the set of colors that appear on the vertices of $S$ under $\alpha$.
Let $P_{n}$, $C_{n}$, and $K_{n}$ denote the path, cycle, and complete graph on $n$ vertices, respectively. Let $K_{p,q}$ denote the complete bipartite graph with $p$ vertices in one set of the bipartition and $q$ vertices in the other. A graph $H$ is said to be a subgraph of $G$, if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. A subgraph of $G$ induced by a subset $S\subseteq V(G)$ is the subgraph of $G$ with vertex-set $S$ and edge-set all edges of $G$ which have both ends in $S$; it is denoted by $G[S]$. An induced subgraph $H$ of $G$ is said to be proper if $V(H)\subset V(G)$. For a subset $S\subseteq V(G)$, we use $G$-$S$ to denote the subgraph of $G$ obtained by deleting the vertices of $S$ from $G$. For a vertex $v$, we use $G$-$v$ instead of $G$-$\{v\}$. For a subset $M\subseteq E(G)$, we use $G$-$M$ to denote the graph obtained by deleting $M$ from $G$, that is, the graph with vertex-set $V(G)$ and edge-set $E(G)\setminus M$.
The complement of a graph $G$, denoted co-$G$, is the graph with the same vertices as $G$ such that two vertices are adjacent in co-$G$ if and only if they are non-adjacent in $G$. A component of a graph $G$ is a maximal connected subgraph of $G$. For two vertex-disjoint graphs $G$ and $H$, the disjoint union of $G$ and $H$, denoted by $G+H$, is the graph with vertex-set $V(G)\cup V(H)$ and edge-set $E(G)\cup E(H)$. For a positive integer $r$, we use $rG$ to denote the graph consisting of the disjoint union of $r$ copies of $G$. The join of two graphs $G_{1}$ and $G_{2}$ is the graph obtained from the disjoint union of $G_{1}$ and $G_{2}$ by adding edges from each vertex of $G_{1}$ to every vertex of $G_{2}$. If $G$ is disconnected, then it is easy to see that $G$ is recolorable if and only if every component of $G$ is recolorable. So we may assume that $G$ is connected when appropriate. We also note that if $G$ is a join of two graphs $G_{1}$ and $G_{2}$, then $G$ is recolorable if both $G_{1}$ and $G_{2}$ are recolorable [1].
3 Modular Decomposition
A module $S$ is said to be maximal if $S\subset V(G)$ and it is not contained in a larger module $M\subset V(G)$. Two sets $X$ and $Y$ overlap if $X$-$Y$, $Y$-$X$, and $X\cap Y$ are all non-empty. If two modules $S_{1}$ and $S_{2}$ of $G$ overlap, then it is easy to see that $S_{1}$-$S_{2}$, $S_{2}$-$S_{1}$, $S_{1}\cap S_{2}$, and $S_{1}\cup S_{2}$ are also modules of $G$. If $G$ is neither a join nor a disjoint union of two graphs, then any two maximal modules of $G$ are disjoint, and hence $V(G)$ can be partitioned into maximal modules $S_{1}$, $S_{2}$,…, $S_{m}$. Partitioning $V(G)$ into maximal modules can be done in linear-time; see [9] for more information.
Let $G$ be a graph which is neither a join nor a disjoint union of two graphs; the skeleton of $G$, denoted $G^{*}$, is the graph obtained by contracting each maximal module to a single vertex (and removing any loops and multiple edges created). It can also be obtained from $G$ by recursively substituting a vertex for each maximal module in $G$. Note that $G^{*}$ is a prime subgraph of $G$, induced by a set containing exactly one vertex from each maximal module in $G$.
For the remainder of this section, we use the following decomposition. Let $G$ be a graph which is neither a join nor a disjoint union of two graphs. Let $S_{1}$,…, $S_{m}$ be a partition of $V(G)$ into maximal modules. Note that $S_{p}\cap S_{q}$ = $\emptyset$ for all $p\neq q$, $p$ and $q\in[m]$. Let $\chi(G)=k$ and let $\chi(G[S_{p}])=k_{p}$, for all $p\in[m]$. For a graph $G$, the clique skeleton of $G$ is the graph obtained from $G$ by recursively substituting a clique $Q_{p}$ of size $k_{p}$ for each $S_{p}$ in $G$, where $p\in[m]$. We use $G(Q_{1},\dots,Q_{m})$ to denote the clique skeleton of $G$. Similar operations have been defined in the literature, for example see [15]. Let $H$ = $G(Q_{1},\dots,Q_{m})$ be the clique skeleton of $G$. Then for any $k$-coloring $\alpha$ of $G$, there exists a $k$-coloring $\beta$ of $H$ such that $\beta(Q_{p})\subseteq\alpha(S_{p})$, for all $p\in[m]$. Note that such a coloring $\beta$ of $H$ can be obtained by coloring each $Q_{p}$ with colors from the set $\alpha(S_{p})$, for all $p\in[m]$. Similarly, given any $k$-coloring $\beta$ of $H$, there exists a $k$-coloring $\alpha$ of $G$ such that for all $p\in[m]$, $\alpha(S_{p})$ = $\beta(Q_{p})$. Thus $\chi(G)$ = $\chi(H)$. Note that the clique skeleton of a graph $G$ is a blowup of the skeleton of $G$.
Lemma 1.
Let $G$ be a graph such that $V(G)$ can be partitioned into maximal modules $S_{1}$,…, $S_{m}$. Assume every proper induced subgraph of $G$ is recolorable. Let $\alpha$ and $\beta$ be two $\chi$-colorings of $G$ such that for each $S_{p}$ of $G$, $\alpha(S_{p})\subseteq\beta(S_{p})$. Then there is a path between $\alpha$ and $\beta$ in $R_{\ell}(G)$, for all $\ell\geq\chi(G)$+1.
Proof.
Let $G$ be a graph such that $V(G)$ can be partitioned into maximal modules $S_{1}$,…, $S_{m}$. Let $\gamma$ be a coloring of $G$ and let $p\in[m]$. For any vertex $v\in S_{p}$, since its neighbors not in $S_{p}$ are complete to $S_{p}$, they will not be colored with a color from the set $\gamma(S_{p})$ under $\gamma$. Let $\ell\geq\chi(G)$+1 and let $\{1,\dots,\ell\}$ be the set of available colors. Let $\alpha$ and $\beta$ be two $\chi$-colorings of $G$ such that $\alpha(S_{p})\subseteq\beta(S_{p})$, for all $p\in[m]$. Then there exists a color, say $c$, that does not appear on $V(G)$ under either $\alpha$ or $\beta$. Since $G[S_{p}]$ is recolorable for all $p\in[m]$, we can use the color $c$ to recolor the vertices in $S_{p}$ from $\alpha$ to $\beta$, for each $p\in[m]$ one at a time, without changing the colors on vertices in $S_{j}$ where $j\neq p$. ∎
Lemma 2.
Let $G$ be a graph such that $V(G)$ can be partitioned into maximal modules $S_{1}$,…, $S_{m}$. Let $H$ = $G(Q_{1},\dots,Q_{m})$ be the clique skeleton of $G$. Let $\alpha$ be any $\chi$-coloring of $G$. Let $\beta_{1}$ and $\beta_{2}$ be any two $\chi$-colorings of $H$ such that for each $S_{p}$ of $G$, $\beta_{1}(Q_{p})\subseteq\alpha(S_{p})$ and $\beta_{2}(Q_{p})\subseteq\alpha(S_{p})$. Then there is a path between $\beta_{1}$ and $\beta_{2}$ in $R_{\ell}(H)$, for all $\ell\geq\chi(H)$+1.
Proof.
Assume the hypotheses. Let $\ell\geq\chi(H)$+1. Let $\{1,\dots,\ell\}$ be the set of available colors. Since $\alpha$ is $\chi$-coloring of $G$, there exists a color, say c, that does not appear on $V(G)$ under $\alpha$. Since $\beta_{1}(Q_{p})\subseteq\alpha(S_{p})$ and $\beta_{2}(Q_{p})\subseteq\alpha(S_{p})$, for all $p\in[m]$, the color $c$ also does not appear on $V(H)$ under either $\beta_{1}$ or $\beta_{2}$. Since each $Q_{p}$ is a clique, it is recolorable. Recolor the vertices in $Q_{p}$ from $\beta_{1}$ to $\beta_{2}$ using the extra color $c$, for each $p\in[m]$ one at a time, without changing the colors on vertices in $Q_{j}$ where $j\neq p$. ∎
Lemma 3.
Let $G$ be a graph such that $V(G)$ can be partitioned into maximal modules $S_{1}$,…, $S_{m}$. Let $H$ = $G(Q_{1},\dots,Q_{m})$ be the clique skeleton of $G$. Let $\ell\geq\chi(G)$+1. Let $\alpha$ be an $\ell$-coloring of $G$ and let $\beta$ be a $\chi$-coloring of $G$. Let $\alpha^{{}^{\prime}}$ be an $\ell$-coloring of $H$ such that for all $p\in[m]$, $\alpha^{{}^{\prime}}(Q_{p})\subseteq\alpha(S_{p})$. Let $\beta^{{}^{\prime}}$ be a $\chi$-coloring of $H$ such that for all $p\in[m]$, $\beta^{{}^{\prime}}(Q_{p})\subseteq\beta(S_{p})$. If there is a path between $\alpha$ and $\beta$ in $R_{\ell}(G)$, then there is a path between $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ in $R_{\ell}(H)$.
Proof.
Assume the hypotheses. Let $\mathbb{P}$ be a path between $\alpha$ and $\beta$ in $R_{\ell}(G)$. Let $\alpha_{i}$ and $\alpha_{i+1}$ be two $\ell$-colorings of $G$, which are adjacent in $\mathbb{P}$. Let $\alpha_{i}^{{}^{\prime}}$ be any $\ell$-coloring of $H$ such that for all $p\in[m]$, $\alpha_{i}^{{}^{\prime}}(Q_{p})\subseteq\alpha_{i}(S_{p})$. We prove that there is a path from $\alpha_{i}^{{}^{\prime}}$ to some $\ell$-coloring of $H$, say $\alpha_{i+1}^{{}^{\prime}}$, in $R_{\ell}(H)$ such that for all $p\in[m]$, $\alpha_{i+1}^{{}^{\prime}}(Q_{p})\subseteq\alpha_{i+1}(S_{p})$.
Since $\alpha_{i}$ and $\alpha_{i+1}$ are adjacent in $R_{\ell}(G)$, they differ on a single vertex, say $v$. Without loss of generality, let $v\in S_{j}$ for some $j\in[m]$. Note that there are no vertices in the closed neighborhood of $v$ colored $\alpha_{i+1}(v)$ under the coloring $\alpha_{i}$ of $G$. Thus, there are no vertices in the closed neighborhood of any vertex in $Q_{j}$ colored $\alpha_{i+1}(v)$ under the coloring $\alpha_{i}^{{}^{\prime}}$ of $H$. Let $\alpha_{i+1}^{{}^{\prime}}$ be the coloring obtained from $\alpha_{i}^{{}^{\prime}}$ by recoloring the vertex in $Q_{j}$ colored $\alpha_{i}(v)$ with the color $\alpha_{i+1}(v)$. It follows that there is a path between $\alpha_{i}^{{}^{\prime}}$ and $\alpha_{i+1}^{{}^{\prime}}$ in $R_{\ell}(H)$ such that for all $p\in[m]$, $\alpha_{i+1}^{{}^{\prime}}(Q_{p})\subseteq\alpha_{i+1}(S_{p})$. Thus, if there exists a path between $\alpha$ and $\beta$ in $R_{\ell}(G)$, then there exists a path from $\alpha^{{}^{\prime}}$ to some coloring $\beta^{*}$ in $R_{\ell}(H)$ such that for all $p\in[m]$, $\beta^{*}(Q_{p})\subseteq\beta(S_{p})$. Since $\beta^{*}$ and $\beta^{{}^{\prime}}$ are two $\chi$-colorings of $H$ such that for all $p\in[m]$, $\beta^{*}(Q_{p})\subseteq\beta(S_{p})$ and $\beta^{{}^{\prime}}(Q_{p})\subseteq\beta(S_{p})$, by Lemma 2 there is a path between $\beta^{*}$ and $\beta^{{}^{\prime}}$ in $R_{\ell}(H)$. Therefore, there is a path between $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ in $R_{\ell}(H)$. ∎
Lemma 4.
Let $G$ be a graph such that $V(G)$ can be partitioned into maximal modules $S_{1}$,…, $S_{m}$. Assume every proper induced subgraph of $G$ is recolorable. Let $H$ = $G(Q_{1},\dots,Q_{m})$ be the clique skeleton of $G$. Let $\ell\geq\chi(G)$+1. Let $\alpha$ and $\beta$ be two $\chi$-colorings of $G$. Let $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ be two $\chi$-colorings of $H$ such that for all $p\in[m]$, $\alpha^{{}^{\prime}}(Q_{p})\subseteq\alpha(S_{p})$ and $\beta^{{}^{\prime}}(Q_{p})\subseteq\beta(S_{p})$. Then there exists a path between $\alpha$ and $\beta$ in $R_{\ell}(G)$ if and only if there exists a path between $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ in $R_{\ell}(H)$.
Proof.
Assume the hypotheses.
First assume that there is a path between $\alpha$ and $\beta$ in $R_{\ell}(G)$. Then by Lemma 3, there is a path between $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ in $R_{\ell}(H)$.
Now assume that there is a path between $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ in $R_{\ell}(H)$. Let $\alpha^{*}$ and $\beta^{*}$ be two $\chi$-colorings of $G$ such that for all $p\in[m]$, $\alpha^{*}(S_{p})$ = $\alpha^{{}^{\prime}}(Q_{p})$ and $\beta^{*}(S_{p})$ = $\beta^{{}^{\prime}}(Q_{p})$. Then since for all $p\in[m]$, $\alpha^{{}^{\prime}}(Q_{p})\subseteq\alpha(S_{p})$ and $\beta^{{}^{\prime}}(Q_{p})\subseteq\beta(S_{p})$, clearly for all $p\in[m]$, $\alpha^{*}(S_{p})\subseteq\alpha(S_{p})$ and $\beta^{*}(S_{p})\subseteq\beta(S_{p})$. Thus by Lemma 1, there is a path from $\alpha$ to $\alpha^{*}$ and there is a path from $\beta$ to $\beta^{*}$ in $R_{\ell}(G)$, for any $\ell\geq\chi(G)$+1. We will now prove that there is a path from $\alpha^{*}$ to $\beta^{*}$ in $R_{\ell}(G)$.
Let $\mathbb{P}$ be a path between $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ in $R_{\ell}(H)$. Let $\alpha_{i}^{{}^{\prime}}$ and $\alpha_{i+1}^{{}^{\prime}}$ be two $\ell$-colorings of $H$ which are adjacent in $\mathbb{P}$. Let $\alpha_{i}$ be an $\ell$-coloring of $G$ such that for all $p\in[m]$, $\alpha_{i}(S_{p})=\alpha_{i}^{{}^{\prime}}(Q_{p})$. We prove that there is a path from $\alpha_{i}$ to some $\ell$-coloring of $G$, say $\alpha_{i+1}$, in $R_{\ell}(G)$, such that $\alpha_{i+1}(S_{p})$ = $\alpha_{i+1}^{{}^{\prime}}(Q_{p})$, for all $p\in[m]$. Since $\alpha_{i}^{{}^{\prime}}$ and $\alpha_{i+1}^{{}^{\prime}}$ are adjacent in $R_{\ell}(H)$, they differ on a single vertex. Let $v\in Q_{j}\subseteq V(H)$, where $\alpha_{i}^{{}^{\prime}}(v)\neq\alpha_{i+1}^{{}^{\prime}}(v)$. Note that there are no vertices in the closed neighborhood of $v$ colored $\alpha_{i+1}^{{}^{\prime}}(v)$ under the coloring $\alpha_{i}^{{}^{\prime}}$ of $H$. Thus there are no vertices in the closed neighborhood of any vertex in $S_{j}$ colored $\alpha_{i+1}^{{}^{\prime}}(v)$ under the coloring $\alpha_{i}$ of $G$. Let $\alpha_{i+1}$ be an $\ell$-coloring of $G$ obtained from $\alpha_{i}$ by recoloring every vertex in $S_{j}$ colored $\alpha_{i}^{{}^{\prime}}(v)$ with the color $\alpha_{i+1}^{{}^{\prime}}(v)$, one at a time. This corresponds to a path between $\alpha_{i}$ and $\alpha_{i+1}$ in $R_{\ell}(G)$, where $\alpha_{i+1}(S_{p})$ = $\alpha_{i+1}^{{}^{\prime}}(Q_{p})$ for all $p\in[m]$. Hence, if there is a path between $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ in $R_{\ell}(H)$, then there is a path from $\alpha^{*}$ to some coloring $\gamma$ in $R_{\ell}(G)$, such that for all $p\in[m]$, $\gamma(S_{p})$ = $\beta^{{}^{\prime}}(Q_{p})$. Since $\beta^{*}$ and $\gamma$ are two $\chi$-colorings of $G$ such that for all $p\in[m]$, $\beta^{*}(S_{p})$ = $\gamma(S_{p})$, by Lemma 1, there is a path between $\beta^{*}$ and $\gamma$ in $R_{\ell}(G)$. Therefore, there is a path between $\alpha^{*}$ and $\beta^{*}$ in $R_{\ell}(G)$. ∎
Theorem 6.
Let $G$ be a graph such that $V(G)$ can be partitioned into maximal modules $S_{1}$,…, $S_{m}$. Let $H$ = $G(Q_{1},\dots,Q_{m})$ be the clique skeleton of $G$. Let $\ell\geq\chi(G)$+1. If every proper induced subgraph of $G$ is recolorable, then $R_{\ell}(G)$ is connected if and only if $R_{\ell}(H)$ is connected.
Proof.
Assume the hypotheses.
We first prove that, for any $\ell\geq\chi(G)$+1, if $R_{\ell}(G)$ is connected, then $R_{\ell}(H)$ is connected. Let $\alpha^{{}^{\prime}}$ be any $\ell$-coloring of $H$ and let $\beta^{{}^{\prime}}$ be a $\chi$-coloring of $H$. Let $\alpha$ be an $\ell$-coloring of $G$ and let $\beta$ be a $\chi$-coloring of $G$, such that for all $p\in[m]$, $\alpha(S_{p})$ = $\alpha^{{}^{\prime}}(Q_{p})$ and $\beta(S_{p})$ = $\beta^{{}^{\prime}}(Q_{p})$. Since $R_{\ell}(G)$ is connected, there is a path between $\alpha$ and $\beta$ in $R_{\ell}(G)$. Then by Lemma 3, there is a path between $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ in $R_{\ell}(H)$. Therefore, $R_{\ell}(H)$ is connected.
We now prove that, for any $\ell\geq\chi(G)$+1, if $R_{\ell}(H)$ is connected, then $R_{\ell}(G)$ is connected. Note that for each $\chi$-coloring $\mu$ of $G$ there is a $\chi$-coloring $\mu^{{}^{\prime}}$ of $H$ such that for all $p\in[m]$, $\mu^{{}^{\prime}}(Q_{p})\subseteq\mu(S_{p})$. Similarly, for each $\chi$-coloring $\mu^{{}^{\prime}}$ of $H$ there is a $\chi$-coloring $\mu$ of $G$ such that for all $p\in[m]$, $\mu(S_{p})$ = $\mu^{{}^{\prime}}(Q_{p})$. Hence, if $R_{\ell}(H)$ is connected, then by Lemma 4, there is a path between any two $\chi$-colorings of $G$ in $R_{\ell}(G)$. Thus, to prove that $R_{\ell}(G)$ is connected, it is sufficient to prove that given any $\ell$-coloring of $G$, we can reach a $\chi$-coloring in $R_{\ell}(G)$.
Let $\alpha$ be an $\ell$-coloring of $G$. Since for all $p\in[m]$, $G[S_{p}]$ is recolorable, recolor the vertices in each $S_{p}$, one at a time, to use at most $\chi(G[S_{p}])$ colors without changing the colors on vertices in $S_{j}$, $j\neq p$, to obtain a coloring $\beta$ of $G$. Let $\beta^{{}^{\prime}}$ be an $\ell$-coloring of $H$ such that for all $p\in[m]$, $\beta^{{}^{\prime}}(Q_{p})=\beta(S_{p})$. Let $\gamma^{{}^{\prime}}$ be a $\chi$-coloring of $H$. Since $R_{\ell}(H)$ is connected, there is a path $\mathbb{P}$ between $\beta^{{}^{\prime}}$ and $\gamma^{{}^{\prime}}$ in $R_{\ell}(H)$. Let $\alpha_{i}^{{}^{\prime}}$ and $\alpha_{i+1}^{{}^{\prime}}$ be two $\ell$-colorings of $H$ which are adjacent in the path $\mathbb{P}$. Let $\alpha_{i}$ be an $\ell$-coloring of $G$ such that for all $p\in[m]$, $\alpha_{i}(S_{p})=\alpha_{i}^{{}^{\prime}}(Q_{p})$. We prove that there is a path from $\alpha_{i}$ to some $\ell$-coloring of $G$, say $\alpha_{i+1}$, in $R_{\ell}(G)$, such that for all $p\in[m]$, $\alpha_{i+1}(S_{p})$ = $\alpha_{i+1}^{{}^{\prime}}(Q_{p})$. Since $\alpha_{i}^{{}^{\prime}}$ and $\alpha_{i+1}^{{}^{\prime}}$ are adjacent in $R_{\ell}(H)$, they differ on a single vertex, say $v$. Let $v\in Q_{j}$, for some $j\in[m]$. Note that there are no vertices in the closed neighborhood of $v$ colored $\alpha_{i+1}^{{}^{\prime}}(v)$ under the coloring $\alpha_{i}^{{}^{\prime}}$ of $H$. Thus there are no vertices in the closed neighborhood of any vertex in $S_{j}$ colored $\alpha_{i+1}^{{}^{\prime}}(v)$ under the coloring $\alpha_{i}$ of $G$. Let $\alpha_{i+1}$ be an $\ell$-coloring of $G$ obtained from $\alpha_{i}$ by recoloring every vertex in $S_{j}$ colored $\alpha_{i}^{{}^{\prime}}(v)$, one at a time, with the color $\alpha_{i+1}^{{}^{\prime}}(v)$. This corresponds to a path between $\alpha_{i}$ and $\alpha_{i+1}$ in $R_{\ell}(G)$, where $\alpha_{i+1}(S_{p})=\alpha_{i+1}^{{}^{\prime}}(Q_{p})$ for all $p\in[m]$. Hence, if there is a path from $\beta^{{}^{\prime}}$ to $\gamma^{{}^{\prime}}$ in $R_{\ell}(H)$, then we can recolor the vertices of $G$ from any $\ell$-coloring $\alpha$ of $G$ to some $\chi$-coloring of $G$. ∎
The results below are well-known, for example see [1, 3].
Lemma 5.
If $G$ is a join of two recolorable graphs $G_{1}$ and $G_{2}$, then $G$ is recolorable.
Lemma 6.
Let $G$ be a graph with non-adjacent vertices $u$ and $v$, such that $N(u)\subseteq N(v)$. Then either
(i)  $G$ is recolorable or
(ii) $G$-$u$ is not recolorable.
We now prove Theorem 1.
Proof of Theorem 1.
Let $\mathcal{G}$ be a hereditary class of graphs and let $\mathcal{H}$ be the set of prime graphs in $\mathcal{G}$. Assume that for all $H\in\mathcal{H}$, every blowup of $H$ is recolorable. If possible, let $G\in\mathcal{G}$ be a vertex-minimal graph that is not recolorable. Then $G$ is neither a join nor a disjoint union of two graphs. Let $S_{1}$,…, $S_{m}$ be a partition of $V(G)$ into maximal modules. Let $G^{*}$ be the skeleton of $G$. Then $G^{*}$ is a prime induced subgraph of $G$ and hence is in $\mathcal{H}$. Since $G^{*}\in\mathcal{H}$, from the hypothesis, every blowup of $G^{*}$ is recolorable. Let $H^{*}$ = $G(Q_{1},\dots,Q_{m})$ be the clique skeleton of $G$. Note that $H^{*}$ is a blowup of $G^{*}$ and hence is recolorable. By the vertex minimality of $G$, every proper induced subgraph of $G$ is recolorable. Therefore, since $H^{*}$ is recolorable, by Theorem 6, $G$ is recolorable, a contradiction. ∎
Let $\mathcal{G}$ be a hereditary class of graphs and let $\mathcal{H}$ be the set of prime graphs in $\mathcal{G}$. It follows from Theorem 1 that to determine if every $G\in\mathcal{G}$ is recolorable, it is sufficient to determine if every blowup of every $H\in\mathcal{H}$ is recolorable. We can further ask: to determine if every $G\in\mathcal{G}$ is recolorable, is it sufficient to determine if every $H\in\mathcal{H}$ is recolorable? We answer this in the positive for $2K_{2}$-free graphs.
Theorem 7.
Let $\mathcal{G}$ be a hereditary class of graphs such that every $G\in\mathcal{G}$ is $K_{2}$-free. If every prime graph in $\mathcal{G}$ is recolorable, then every graph in $\mathcal{G}$ is recolorable.
Proof.
Assume the hypothesis. If possible, let $G\in\mathcal{G}$ be a vertex-minimal graph that is not recolorable. Then $G$ is not prime. Since $G$ is not prime, it contains a maximal module with at least two vertices, say $S$. Let $u,v\in S$. If $S$ is an independent set, then $N(u)$ = $N(v)$. Then by Lemma 6, $G$ is not vertex-minimal, a contradiction. So $S$ must contain an edge. Without loss of generality, let $uv\in E(G)$. If $S$ is complete to $V(G)\setminus S$, then $G$ is a join of $G[S]$ and $G$-$S$. Then by Lemma 5, $G$ is not vertex-minimal, a contradiction. Hence there is a vertex $x\in V(G)\setminus S$ that is non-adjacent to some vertex in $S$. Since $S$ is a module, $x$ is non-adjacent to every vertex in $S$. If $N(x)\subseteq N(v)$, then by Lemma 6, $G$ is not vertex-minimal, a contradiction. Thus there is a vertex $y\in N(x)\setminus N(v)$. Then $y\in V(G)\setminus S$ since $x$ is non-adjacent to every vertex in $S$. Since $S$ is a module and $y$ is non-adjacent to $v\in S$, $y$ is non-adjacent to every vertex in $S$. Then $\{u,v,x,y\}$ induces a $K_{2}$, a contradiction. ∎
Lemma 7.
([1]) Let $G$ contain a vertex $v$ of degree at most $\chi(G)$-1. Then, either
(i)  $G$ is recolorable or
(ii) $G$-$v$ is not recolorable.
Theorem 8.
Let $\mathcal{G}$ be a hereditary class of graphs such that every $G\in\mathcal{G}$ is diamond-free. If every prime graph in $\mathcal{G}$ is recolorable, then every graph in $\mathcal{G}$ is recolorable.
Proof.
Assume the hypothesis. If possible, let $G\in\mathcal{G}$ be a vertex-minimal graph that is not recolorable. Then $G$ is connected and not prime. Since $G$ is not prime, it contains a non-trivial module, say $S$, with at least two vertices.
Since $G$ is connected, there exists a vertex $z\in V(G)\setminus S$ adjacent to every vertex in $S$. If $G[S]$ contains an induced $P_{3}$, induced by, say, the set $P\subseteq S$, then $P\cup\{z\}$ induces a diamond in $G$, a contradiction. So $G[S]$ is $P_{3}$-free and hence $G[S]$ is a disjoint union of cliques.
If there are adjacent vertices $u$ and $v$ in $S$, then we claim that $N[u]=N[v]$ is a clique. Since $S$ is a module and a disjoint union of cliques, $N[u]=N[v]$. If $N[u]$ is not a clique, then there exist two non-adjacent vertices $x$ and $y$ in $V(G)\setminus S$ adjacent to both $u$ and $v$, a contradiction to the fact that $G$ is diamond-free. Thus $N[u]=N[v]$ is a clique. Therefore, where $d(u)$ denotes the degree of the vertex $u$ in $G$, $d(u)\leq\omega(G)$-1 $\leq\chi(G)$-1. By vertex minimality of $G$, the graph $G$-$u$ must be recolorable. Then by Lemma 7, $G$ is recolorable, a contradiction.
If there are no adjacent vertices in $S$, then $S$ is an independent set in $G$. Let $\{u,v\}\subseteq S$. Since $S$ is a module, $N(u)$ = $N(v)$. By vertex minimality of $G$, the graph $G$-$u$ must be recolorable. Then by Lemma 6, $G$ is recolorable, a contradiction. ∎
4 $P_{5}$-free graphs
In this section, we use the results from previous section to prove recolorability results for several subclasses of $P_{5}$-free graphs. A clique cutset $Q$ of a graph $G$ is a clique in $G$ such that $G$-$Q$ has more components than $G$. A clique cutset $Q$ of a graph $G$ is called a tight clique cutset if there exists a component $H$ of $G$-$Q$ which is complete to $Q$.
We use the following results in our proof.
Lemma 8.
([2]) Suppose $G$ has a tight clique cutset. If every proper induced subgraph of $G$ is recolorable, then $G$ is recolorable.
Lemma 9.
[…]
Proof.
[…]
A graph is said to be chordal if it does not contain an induced cycle of length more than three.
Lemma 10.
Every blowup of a chordal graph is recolorable.
Proof.
Clearly, every blowup of a chordal graph is chordal. Thus the lemma follows from the result proved in [5] that every chordal graph is recolorable. ∎
Lemma 11.
Every blowup of a prime graph with at most 5 vertices is recolorable.
Proof.
Let $G$ be any prime graph with at most 5 vertices. If $G$ is ($P_{5}$, $C_{4}$)-free, then every blowup of $G$ is ($P_{5}$, $C_{4}$)-free. Thus the lemma follows from the result proved in [2] that every ($P_{5}$, $C_{4}$)-free graph is recolorable. Note that there are no prime graphs with at most 4 vertices containing an induced $P_{5}$ or an induced $C_{4}$.
[…]
Two edges in a graph are said to be independent if they do not share an end vertex. A matching is a set of mutually independent edges. A matched co-bipartite graph is the complement of $K_{p,q}$-$M$, where $M$ is a maximum matching, for some $p,q\geq 1$. Note that every matched co-bipartite graph is a co-bipartite graph. We use the modular decomposition of ($P_{5}$, diamond)-free graphs given by Brandstädt [7] to prove that every ($P_{5}$, diamond)-free graph is recolorable.
Theorem 9.
([7]) If $G$ is a prime ($P_{5}$, diamond)-free graph containing a $2K_{2}$, then $G$ is a matched co-bipartite graph.
Theorem 10.
([1]) Every ($K_{2}$, diamond)-free graph is recolorable.
We now have all the ingredients to prove Theorem 2.
Proof of Theorem 2.
[…]
Lemma 12.
([14]) Let $G$ = ($V$, $E$) be a prime graph on $n$ vertices. Then $G$ is bipartite and $P_{5}$-free if and only if the following conditions hold:
  • There exists a partition of $V(G)$ into two independent sets $B$ = $\{b_{1},b_{2},\dots,b_{\frac{n}{2}}\}$ and $W$ = $\{w_{1},w_{2},\dots,w_{\frac{n}{2}}\}$.
  • The neighbors of $b_{i}$ ($i$ = 1,$\dots,\frac{n}{2}$) are precisely $w_{1},\dots,w_{\frac{n}{2}-i+1}$.
Lemma 13.
Every blowup of a prime $P_{5}$-free bipartite graph is recolorable.
Proof.
If possible, let $H$ be a vertex-minimal minimal counter example, say $H$ is a blowup of some prime $P_{5}$-free bipartite graph $G$. By Lemma 12, $G$ has a vertex of degree 1. The neighbor of this degree 1 vertex is a tight clique cutset of $G$. Thus $H$ has a tight clique cutset. Every proper induced subgraph of $H$ is also a blowup of some prime $P_{5}$-free bipartite graph and hence recolorable. Then by Lemma 8, $H$ is recolorable, a contradiction. ∎
Lemma 14.
Every blowup of a co-bipartite graph is recolorable.
Proof.
[…]
Theorem 11.
([12]) If $G$ is a prime ($P_{5}$, house, bull)-free graph on 6 or more vertices without a universal vertex, then $G$ is either a bipartite graph or a co-bipartite graph.
We now prove Theorem 3.
Proof of Theorem 3.
By Theorem 1, it is sufficient to prove that every blowup of every prime ($P_{5}$, house, bull)-free graph is recolorable. Let $G$ be a prime ($P_{5}$, house, bull)-free graph. By Lemma 5, we may assume that $G$ does not have a universal vertex. By Lemma 11, we may assume that $G$ has at least 6 vertices. By Theorem 11, $G$ is a bipartite graph or a co-bipartite graph. Then the result follows from Lemma 13 and Lemma 14. ∎
A thin spider is a graph whose vertex set can be decomposed into a clique $K$ and an independent set $S$, such that $|K|$ = $|S|$ or $|K|$ = $|S|$+1, and the edges between $S$ and $K$ form a matching, leaving at most one vertex of $K$ with no neighbor in $S$.
Lemma 15.
Every blowup of a thin spider or the complement of a thin spider is recolorable.
Proof.
Let $H$ be a blowup of a thin spider or the complement of a thin spider. Since every thin spider and every complement of a thin spider is chordal, the graph $H$ is recolorable by Lemma 10. ∎
Theorem 12.
([13]) Let $G$ be a prime semi-$P_{4}$-sparse graph. Then either $G$ or its complement is a bipartite graph or a thin spider.
We now prove Theorem 4.
Proof of Theorem 4.
By Theorem 1, it is sufficient to prove that every blowup of every prime semi-$P_{4}$-sparse graph is recolorable. Let $G$ be a prime semi-$P_{4}$-sparse graph. Note that $G$ is $P_{5}$-free. By Theorem 12, $G$ or its complement is a bipartite graph or a thin spider. If $G$ or its complement is a bipartite graph, then the result follows from Lemma 13 or Lemma 14, respectively. If $G$ or its complement is a thin spider, then the result follows from Lemma 15. ∎
Theorem 13.
([14]) Every prime ($P_{5}$, house)-free graph is either a $C_{5}$ or $C_{5}$-free.
We now prove Theorem 5.
Proof of Theorem 5.
[…]
5 Conclusion
Let $\mathcal{G}$ be a hereditary class of graphs. In Theorem 1, we proved that every graph in $\mathcal{G}$ is recolorable if every blowup of every prime graph in $\mathcal{G}$ is recolorable. We have two questions.
1351351246246
Figure 2: A graph $G$ (left) and a blowup $G^{{}^{\prime}}$ of it (right).
Question 1: If every prime graph in $\mathcal{G}$ is recolorable, then is every graph in $\mathcal{G}$ recolorable?
We answered Question 1 positively for $K_{2}$-free graphs and would like to know if it is true in general.
Furthermore, we can ask if every blowup of a prime recolorable graph recolorable. We answer this in the negative.
It is easy to see that the graph $G$ in Figure 2 (left) is a prime 3-chromatic graph and it is recolorable, but the blowup $G^{{}^{\prime}}$ of $G$ (right) is 5-chromatic and $R_{6}(G^{{}^{\prime}})$ is disconnected and hence $G^{{}^{\prime}}$ is not recolorable. However, we can see that $G$ has a prime induced subgraph, $C_{6}$, which is not recolorable.
We conjecture the following:
Conjecture 1.
Let $G$ be a prime graph. Every blowup of $G$ is recolorable if and only if every induced subgraph of $G$ is recolorable.
Acknowledgement
We thank Chính T. Hoàng for helpful suggestions.

### v2

@@
 A graph $G$ is said to be recolorable if $R_{\ell}(G)$ is connected for all $\ell\geq\chi(G)$+1.
-We use the modular decomposition of several graph classes to prove that the graphs in the class are recolorable.
+We demonstrate how to use the modular decomposition of a graph class to prove that the graphs in the class are recolorable.
 In particular, we prove that every ($P_{5}$, diamond)-free graph, every ($P_{5}$, house, bull)-free graph, and every ($P_{5}$, $C_{5}$, co-fork)-free graph is recolorable.
+A graph is prime if it cannot be decomposed by modular decomposition except into single vertices.
+For a prime graph $H$, we study the complexity of deciding if $H$ is $k$-colorable and the complexity of deciding if there exists a path between two given $k$-colorings in $R_{k}(H)$.
+Suppose $\mathcal{G}$ is a hereditary class of graphs.
+We prove that if every blowup of every prime graph in $\mathcal{G}$ is recolorable, then every graph in $\mathcal{G}$ is recolorable.
+Keywords: reconfiguration graph, $k$-mixing, modular decomposition, $P_{5}$-free graphs.
 1 Introduction
@@
 In this paper we focus on the connectivity of the reconfiguration graph of the $k$-colorings.
-We say a graph $G$ is recolorable if $R_{\ell}(G)$ is connected for all $\ell\geq\chi(G)$+1.
+A graph is said to be $k$-mixing if $R_{k}(G)$ is connected.
+We say a graph $G$ is recolorable if $G$ is $\ell$-mixing for all $\ell\geq\chi(G)$+1.
 Deciding whether there exists a path between two colorings in $R_{k}(G)$ is PSPACE-complete for $k>3$ [7], and can be decided in polynomial time for $k\leq 3$ [9].
@@
 This led researchers to study the problem for various restricted graph classes.
-In [5], a graph class known as $k$-color-dense graphs was introduced and it was proved that every $k$-color-dense graph is recolorable.
+In [6], Bonamy et.
+al.
+introduced the class of $k$-color-dense graphs and proved that every $k$-color-dense graph is recolorable.
 The class of $k$-color-dense graphs includes the class of $k$-colorable chordal graphs.
@@
 For a collection of graphs $\mathcal{H}$, $G$ is $\mathcal{H}$-free if $G$ is $H$-free for every $H\in\mathcal{H}$.
+For more results about reconfiguration of vertex colorings in forbidden induced subgraph classes see [1].
 Bonamy and Bousquet [5] used the tree decomposition of graphs to show that every graph $G$ is ($tw(G)$+2)-mixing, where $tw(G)$ denotes the treewidth of the graph $G$.
-In this paper we use the modular decomposition of a graph $G$ to prove that $G$ is recolorable.
-A set $S\subseteq V(G)$ is a module if $1\leq|S|\leq|V(G)|$ and every vertex in $V(G)\setminus S$ is either adjacent to every vertex of $S$ or no vertex of $S$.
-A module $S$ is said to be non-trivial if $S$ contains at least two vertices.
+In this paper we demonstrate how to use the modular decomposition of a graph $G$ to prove that $G$ is $\ell$-mixing for $\ell\geq\chi(G)$+1.
+A non-empty set $S\subseteq V(G)$ is a module if every vertex in $V(G)\setminus S$ is either adjacent to every vertex of $S$ or no vertex of $S$.
+A module $S\subset V(G)$ is said to be non-trivial if $S$ contains at least two vertices.
 Note that for each $v\in V(G)$, $\{v\}$ is a trivial module in $G$.
@@
 Note that for a collection of graphs $\mathcal{H}$, the class of $\mathcal{H}$-free graphs is hereditary.
-In Section 3, we prove the following.
+In Section 3, for a prime graph $H$, we study the complexity of deciding if $H$ is $k$-colorable and the complexity of deciding if there exists a path between two given $k$-colorings in $R_{k}(H)$.
 Theorem 1.
-Let $\mathcal{G}$ be a hereditary class of graphs and let $\mathcal{H}$ be the set of prime graphs in $\mathcal{G}$.
-If for all $H\in\mathcal{H}$, every blowup of $H$ is recolorable, then for all $G\in\mathcal{G}$, $G$ is recolorable.
-We use the above result in Section 4 along with the modular decomposition of several subclasses of $P_{5}$-free graphs to prove the following.
+Given a prime graph $H$, for all $k\geq 3$, deciding whether $H$ is $k$-colorable is NP-complete.
+Theorem 2.
+Given a prime graph $H$, for all $k\geq 4$, deciding whether there exists a path between two given colorings in $R_{k}(H)$ is PSPACE-complete.
+In Section 4, we will demonstrate how to use modular decomposition in reconfiguration of vertex coloring.
+Theorem 3.
+Let $\mathcal{G}$ be a hereditary class of graphs. If every blowup of every prime graph in $\mathcal{G}$ is recolorable, then every graph in $\mathcal{G}$ is recolorable.
+We use the above result in Section 5 along with the modular decomposition of several subclasses of $P_{5}$-free graphs to prove the following. (See Figure 1 for graphs mentioned in Theorems 4 and 5.)
 Theorem 4.
@@
 We also note that if $G$ is a join of two graphs $G_{1}$ and $G_{2}$, then $G$ is recolorable if both $G_{1}$ and $G_{2}$ are recolorable [2].
-3 Modular Decomposition
+3 Complexity
 A module $S$ is said to be maximal if $S\subset V(G)$ and it is not contained in a larger module $M\subset V(G)$.
@@
 Partitioning $V(G)$ into maximal modules can be done in linear-time; see [10] for more information.
+For a graph $G$, the sibling of $G$ is the graph obtained by adding, for each $x\in V(G)$, a vertex $x^{{}^{\prime}}$ adjacent to only $x$.
+Lemma 1.
+For any connected graph $G$ the sibling of $G$ is prime.
+Proof.
+Let $H$ be the sibling of a connected graph $G$ obtained by adding, for each $x\in V(G)$, a vertex $x^{{}^{\prime}}$ adjacent to only $x$.
+Assume $H$ is not prime.
+Then $H$ contains a non-trivial module $S^{{}^{\prime}}$.
+Suppose $S^{{}^{\prime}}\cap V(G)$ = $\emptyset$.
+Let $x^{{}^{\prime}}$ and $y^{{}^{\prime}}$ be in $S^{{}^{\prime}}$.
+Since $x\in N_{H}(x^{{}^{\prime}})\setminus N_{H}(y^{{}^{\prime}})$, $x$ must be in $S^{{}^{\prime}}$, a contradiction.
+Thus $S^{{}^{\prime}}\cap V(G)\neq\emptyset$.
+Claim 1: For any $x\in S^{{}^{\prime}}\cap V(G)$, $x^{{}^{\prime}}\in S^{{}^{\prime}}$.
+Let $x\in S^{{}^{\prime}}\cap V(G)$.
+Suppose $x^{{}^{\prime}}\notin S^{{}^{\prime}}$.
+Since $S^{{}^{\prime}}$ is a non-trivial module, it must contain some vertex $y$ of $H$ different from $x$.
+Then $x^{{}^{\prime}}$ is adjacent to $x$ but not to $y$, this contradicts that $S^{{}^{\prime}}$ is a module.
+This proves Claim 1.
+Claim 2: For any $x\in S^{{}^{\prime}}\cap V(G)$, $N_{G}(x)\subseteq S^{{}^{\prime}}$.
+Let $x\in S^{{}^{\prime}}\cap V(G)$.
+By Claim 1, $x^{{}^{\prime}}\in S^{{}^{\prime}}\cap V(H)$.
+Since every vertex in $N_{G}(x)$ is adjacent to $x$ but not to $x^{{}^{\prime}}$, we have $N_{G}(x)\subseteq S^{{}^{\prime}}$.
+This proves Claim 2.
+Since $H$ and $G$ are connected and $V(H)=V(G)\cup\{x^{{}^{\prime}}\mid x\in V(G)\}$, by Claims 1 and 2, we have $S^{{}^{\prime}}=V(H)$.
+This contradicts that $S^{{}^{\prime}}$ is a non-trivial module.
+∎
+Theorem 8 ([16]).
+For all $k\geq 3$, deciding whether a connected graph is $k$-colorable is NP-complete.
+We are now ready to prove Theorem 1.
+Proof of Theorem 1.
+Clearly the problem is in NP.
+To prove that it is NP-complete we will give a reduction from a connected graph to a prime graph.
+Let $G$ be a connected graph and let $H$ be the sibling of $G$ obtained by adding, for each $x\in V(G)$, a vertex $x^{{}^{\prime}}$ adjacent to only $x$.
+Note that, for all $k\geq 3$, $H$ is $k$-colorable if and only if $G$ is $k$-colorable.
+By Lemma 1, $H$ is a prime graph.
+This proves the theorem.
+∎
+Theorem 9 ([7]).
+Given a connected graph $G$, for all $k\geq 4$, deciding whether there exists a path between two colorings in $R_{k}(G)$ is PSPACE-complete.
+We will now prove Theorem 2.
+Proof of Theorem 2.
+Clearly the problem is in PSPACE.
+To prove that it is PSPACE-complete we will give a reduction from a connected graph to a prime graph.
+Let $G$ be a connected graph and let $H$ be the sibling of $G$ obtained by adding, for each $x\in V(G)$, a vertex $x^{{}^{\prime}}$ adjacent to only $x$.
+Then, by Lemma 1, $H$ is a prime graph.
+Let $k\geq 4$.
+Let $\alpha$ and $\beta$ be any two $k$-colorings of $G$.
+Let $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ be two $k$-colorings of $H$ obtained by extending $\alpha$ and $\beta$, respectively, by coloring each vertex $x^{{}^{\prime}}\in V(H)\setminus V(G)$ a color different from the color of $x$.
+Since $G$ is an induced subgraph of $H$, any $k$-coloring of $H$ restricted to $V(G)$ is a $k$-coloring of $G$.
+Thus if there is a path between $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ in $R_{k}(H)$, then there is a path between $\alpha$ and $\beta$ in $R_{k}(G)$.
+Assume that there is a path $P$ between $\alpha$ and $\beta$ in $R_{k}(G)$.
+Let $\gamma$ and $\psi$ be any two $k$-colorings of $G$ adjacent in the path $P$.
+Let $\gamma^{{}^{\prime}}$ and $\psi^{{}^{\prime}}$ be any two $k$-colorings of $H$ whose restriction to $V(G)$ is $\gamma$ and $\psi$, respectively.
+We will prove that there is a path between $\gamma^{{}^{\prime}}$ and $\psi{{}^{\prime}}$ in $R_{k}(H)$.
+Since $\gamma$ and $\psi$ are adjacent in the path $P$, there exists a unique vertex $y\in V(G)$ such that $\gamma(y)\neq\psi(y)$.
+Let $y^{{}^{\prime}}\in V(H)\setminus V(G)$ be adjacent to only $y$.
+Starting from $\gamma^{{}^{\prime}}$, recolor $y^{{}^{\prime}}$ with a color not in $\{\gamma(y),\psi(y)\}$.
+Recolor the vertex $y$ with color $\psi(y)$.
+Recolor each vertex $x^{{}^{\prime}}$ in $V(G)\setminus V(H)$, one at a time, with color $\psi^{{}^{\prime}}(x^{{}^{\prime}})$ to obtain the coloring $\psi^{{}^{\prime}}$.
+Hence there is a path between $\alpha$ and $\beta$ in $R_{k}(G)$ if and only if there is a path between $\alpha^{{}^{\prime}}$ and $\beta^{{}^{\prime}}$ in $R_{k}(H)$. This completes the proof. ∎
+4 Recoloring and Modular Decomposition
 Let $G$ be a graph which is neither a join nor a disjoint union of two graphs; the skeleton of $G$, denoted $G^{*}$, is the graph obtained by contracting each maximal module to a single vertex (and removing any loops and multiple edges created).
@@
 Similar operations have been defined in the literature, for example see [17].
+In Figure 2, for a graph $G$, we have illustrated its skeleton and its clique skeleton.
 Let $H$ = $G(Q_{1},\dots,Q_{m})$ be the clique skeleton of $G$.
@@
 Note that the clique skeleton of a graph $G$ is a blowup of the skeleton of $G$.
+$a_{1}$$a_{2}$$b_{1}$$b_{2}$$c_{1}$$d_{1}$$e_{1}$$e_{2}$$e_{3}$$e_{4}$$e_{5}$
+(a) A graph $G$
+$A$$B$$C$$D$$E$
+(b) The skeleton of $G$
+[…]
+(c) The clique skeleton of $G$
+Figure 2: A graph $G$, its skeleton, and its clique skeleton.
 Lemma 2.
@@
 If possible, let $G\in\mathcal{G}$ be a vertex-minimal graph that is not recolorable.
-Then $G$ is neither a join nor a disjoint union of two graphs.
+Then, by Lemma 6, $G$ is neither a join nor a disjoint union of two graphs.
 Let $S_{1}$,…, $S_{m}$ be a partition of $V(G)$ into maximal modules.
@@
 ∎
-Let $\mathcal{G}$ be a hereditary class of graphs and let $\mathcal{H}$ be the set of prime graphs in $\mathcal{G}$.
-It follows from Theorem 1 that to determine if every $G\in\mathcal{G}$ is recolorable, it is sufficient to determine if every blowup of every $H\in\mathcal{H}$ is recolorable.
-We can further ask: to determine if every $G\in\mathcal{G}$ is recolorable, is it sufficient to determine if every $H\in\mathcal{H}$ is recolorable?
-We answer this in the positive for $2K_{2}$-free graphs.
+Let $\mathcal{G}$ be a hereditary class of graphs.
+Then Theorem 3 states that to determine if every graph in $\mathcal{G}$ is recolorable, it is sufficient to determine if every blowup of every prime graph in $\mathcal{G}$ is recolorable.
+We can further ask: To determine if every graph in $\mathcal{G}$ is recolorable, is it sufficient to determine if every prime graph in $\mathcal{G}$ is recolorable?
+We answer this in the positive for $2K_{2}$-free graphs and diamond-free graphs.
 Theorem 11.
@@
 ∎
-Lemma 7.
-([1]) Let $G$ contain a vertex $v$ of degree at most $\chi(G)$-1. Then, either
+Lemma 8 ([2]).
+Let $G$ contain a vertex $v$ of degree at most $\chi(G)$-1. Then, either
 (i)  $G$ is recolorable or
@@
 We use the following results in our proof.
-Lemma 8.
-([2]) Suppose $G$ has a tight clique cutset. If every proper induced subgraph of $G$ is recolorable, then $G$ is recolorable.
+Lemma 9 ([3]).
+Suppose $G$ has a tight clique cutset. If every proper induced subgraph of $G$ is recolorable, then $G$ is recolorable.
 Lemma 10.
@@
 We use the modular decomposition of ($P_{5}$, diamond)-free graphs given by Brandstädt [8] to prove that every ($P_{5}$, diamond)-free graph is recolorable.
-Theorem 9.
-([7]) If $G$ is a prime ($P_{5}$, diamond)-free graph containing a $2K_{2}$, then $G$ is a matched co-bipartite graph.
-Theorem 10.
-([1]) Every ($K_{2}$, diamond)-free graph is recolorable.
+Theorem 13 ([8]).
+If $G$ is a prime ($P_{5}$, diamond)-free graph containing a $2K_{2}$, then $G$ is a matched co-bipartite graph.
+Theorem 14 ([2]).
+Every ($K_{2}$, diamond)-free graph is recolorable.
 We now have all the ingredients to prove Theorem 4.
@@
 […]
-Lemma 12.
-([14]) Let $G$ = ($V$, $E$) be a prime graph on $n$ vertices. Then $G$ is bipartite and $P_{5}$-free if and only if the following conditions hold:
+Lemma 13 ([15]).
+Let $G$ = ($V$, $E$) be a prime graph on $n$ vertices. Then $G$ is bipartite and $P_{5}$-free if and only if the following conditions hold:
   • There exists a partition of $V(G)$ into two independent sets $B$ = $\{b_{1},b_{2},\dots,b_{\frac{n}{2}}\}$ and $W$ = $\{w_{1},w_{2},\dots,w_{\frac{n}{2}}\}$.
@@
 […]
-Theorem 11.
-([12]) If $G$ is a prime ($P_{5}$, house, bull)-free graph on 6 or more vertices without a universal vertex, then $G$ is either a bipartite graph or a co-bipartite graph.
+Theorem 15 ([13]).
+If $G$ is a prime ($P_{5}$, house, bull)-free graph on 6 or more vertices without a universal vertex, then $G$ is either a bipartite graph or a co-bipartite graph.
 We now prove Theorem 5.
@@
 Let $H$ be a blowup of a thin spider or the complement of a thin spider. Since every thin spider and every complement of a thin spider is chordal, the graph $H$ is recolorable by Lemma 11. ∎
-Theorem 12.
-([13]) Let $G$ be a prime semi-$P_{4}$-sparse graph. Then either $G$ or its complement is a bipartite graph or a thin spider.
+Theorem 16 ([14]).
+Let $G$ be a prime semi-$P_{4}$-sparse graph. Then either $G$ or its complement is a bipartite graph or a thin spider.
 We now prove Theorem 6.
@@
 ∎
-Theorem 13.
-([14]) Every prime ($P_{5}$, house)-free graph is either a $C_{5}$ or $C_{5}$-free.
+Theorem 17 ([15]).
+Every prime ($P_{5}$, house)-free graph is either a $C_{5}$ or $C_{5}$-free.
 We now prove Theorem 7.
@@
 Question 1: If every prime graph in $\mathcal{G}$ is recolorable, then is every graph in $\mathcal{G}$ recolorable?
-We answered Question 1 positively for $K_{2}$-free graphs and would like to know if it is true in general.
-Furthermore, we can ask if every blowup of a prime recolorable graph recolorable. We answer this in the negative.
+We answered Question 1 positively for $K_{2}$-free graphs and diamond-free graphs, and would like to know if it is true in general.
+In light of Theorem 3, to answer Question 1 positively, one might check if every blowup of a prime recolorable graph is recolorable. However, this is not true.
 It is easy to see that the graph $G$ in Figure 3 (left) is a prime 3-chromatic graph and it is recolorable, but the blowup $G^{{}^{\prime}}$ of $G$ (right) is 5-chromatic and $R_{6}(G^{{}^{\prime}})$ is disconnected and hence $G^{{}^{\prime}}$ is not recolorable.
@@
 Acknowledgement
-We thank Chính T. Hoàng for helpful suggestions.
+We thank Chính T. Hoàng for helpful suggestions. Some of the work in this paper appeared in [1].